# Shane Cooper and Ilia Kamotski

### v1

@@ 6 Appendix
$(u,\tilde{u})_{H^{1}_{\theta}(\square;\mathbb{C}^{2})}=\left\{\begin{array}[]{lcc}\int_{\square}\nabla u:\overline{\nabla\tilde{u}}&&\theta\neq 0,\\[1.99997pt] \int_{\square}\nabla u:\overline{\nabla\tilde{u}}+\left|\int_{\square}u\right|\left|\int_{\square}\tilde{u}\right|&&\theta=0,\end{array}\right.$ (6.1)
@@
[…]
@@
for arbitrary functions $c_{1},c_{2}$ supported in \$\𝑅\overline{\square\backslash R}$.
@@
$\quad\forall\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}),$
@@
\$\int_{\square\backslash R}c_{1}\overline{\mathrm{div}\,{\phi}}+c_{2}\overline{\mathrm{curl}\,{\phi}}+\int_{\square\backslash R}v_{2,1}\overline{\left(\frac{\phi_{2,1}+\phi_{1,2}}{2}\right)}+v_{2,2}\overline{\left(\frac{\phi_{2,2}-\phi_{1,1}}{2}\right)}$
@@
$\forall\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2})$. Decomposing $\phi$ with respect to the decomposition $H^{1}_{\theta}(\square;\mathbb{C}^{2})=V(\theta)\oplus W(\theta)$, utilising the fact $v$ is an eigenfunction of $\mathcal{B}(\theta)$ with eigenvalue $\lambda=\omega^{2}$ (cf. (2.21) and in particular (2.8)) and utilising Proposition 6.1, it follows that: (6.3) holds if, and only if, the pair$(c_{1},c_{2})$ satisfies
\$\int_{\square\backslash R}c_{1}$ $\overline{\mathrm{div}\,{\tilde{w}}}+c_{2}\overline{\mathrm{curl}\,{\tilde{w}}}=\langle F,\tilde{w}\rangle,\quad\forall\tilde{w}\in W(\theta),$ (6.4)
@@
$\langle F,\tilde{w}\rangle:=\omega^{2}\Big{(}\int_{\square\backslash R}v\cdot\overline{\tilde{w}}$ $+\gamma\int_{R}v_{2}\overline{\tilde{w}_{2}}\Big{)}$
\$-\int_{\square\backslash R}v_{2,1}\overline{\left(\frac{\tilde{w}_{2,1}+\tilde{w}_{1,2}}{2}\right)}+v_{2,2}\overline{\left(\frac{\tilde{w}_{2,2}-\tilde{w}_{1,1}}{2}\right)}-\int_{R}\nabla v_{2}\cdot\overline{\nabla\tilde{w}_{2}}.$
@@
\$\int_{\square\backslash R}\mathrm{div}\,{w}\,\overline{\mathrm{div}\,\tilde{w}}+\mathrm{curl}\,{w}\,\overline{\mathrm{curl}\,{w}}=\langle F,\tilde{w}\rangle\quad\forall\tilde{w}\in W(\theta).$ (6.5)
@@
Here, we shall prove Lemma 3.1. For this, we shall rely on the techniques introduced in [5, Section 3].
For fixed $10<\varepsilon<$ 1, $\theta\in\square^{*}$ and $F\in H^{-1}(\square;\mathbb{C}^{2})$, recall $U_{\varepsilon,\theta}$ is the solution to
$\varepsilon^{-1}a(U_{\varepsilon,\theta},\phi)+b_{\varepsilon}(U_{\varepsilon,\theta},\phi)+c(U_{\varepsilon,\theta},\phi)=\langle F,\phi\rangle,\quad\forall\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}),$ (6.6)
@@
$b(V_{\theta},\tilde{V})=\langle F,\tilde{V}\rangle,\quad\forall\,\tilde{V}\in V(\theta).$
@@
$\varepsilon^{-1}a(W_{\varepsilon,\theta},\phi)+b(W_{\varepsilon,\theta},\phi)+c(W_{\varepsilon,\theta},\phi)=\langle F,\phi\rangle,\quad\forall\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}),$ (6.7)
@@
$\nu_{1}\|\nabla\phi\|^{2}_{L^{2}(\square;\mathbb{C}^{2})}\leq a[\phi]+b[\phi]\leq\nu_{2}\|\nabla\phi\|^{2}_{L^{2}(\square;\mathbb{C}^{2})},\quad\forall\phi\in H^{1}(\square;\mathbb{C}^{2}),$ (6.8)
@@
$\left|b_{\varepsilon}(\phi,\psi)-b(\phi,\psi)\right|\leq\varepsilon L_{b}\sqrt{\varepsilon^{-1}a[\phi]+b_{\varepsilon}[\phi]}\sqrt{a[\psi]+b[\psi]}\quad\forall\phi,\psi\in H^{1}(\square;\mathbb{C}^{2}),$ (6.9)
and for all $10<\varepsilon<$ 1.
@@
for some $\nu_{*}$ independent of $\varepsilon,\theta$ and $F$. To do this, we shall utilise [5, Theorem 3.1]; setting,
$a_{\theta}(\phi,\psi):=a(e^{\mathrm{i}\theta\cdot y}\phi,e^{\mathrm{i}\theta\cdot y}\psi),\qquad\text{and}\quad\ b_{\theta}(\phi,\psi):=b(e^{\mathrm{i}\theta\cdot y}\phi,e^{\mathrm{i}\theta\cdot y}\psi)+c(\phi,\psi),$
@@
$\varepsilon^{-1}a_{\theta}(u_{\varepsilon,\theta},\phi)+b_{\theta}(u_{\varepsilon,\theta},\phi)=\langle f,\phi\rangle,\quad\forall\phi\in H^{1}_{per}(\square;\mathbb{C}^{2});$
@@
$b_{\theta}(v_{\theta},\tilde{v})=\langle f,\tilde{v}\rangle,\quad\forall\tilde{v}\in\mathcal{V}(\theta).$
@@
$\R\mathcal{V}(\theta):=\{v\in H^{1}_{per}(\square;\mathbb{C}^{2})\,|\,\mathrm{div}\,{v}+\mathrm{i}\theta\cdot v=0,\text{ and }\mathrm{curl}\,{v}+\mathrm{i}\theta^{\perp}\cdot v=0\text{ in $\square\backslash R$}\}.$
We now argue as in [5, Section 3]. First note that the closed subspace $\mathcal{V}(\theta)$ is the kernel of the form $a_{\theta}$, precisely
@@
Furthermore, it can be seen that $a_{\theta}$ is coercive on $(\mathcal{V}(\theta))^{\perp}$, the orthogonal complement of $\mathcal{V}(\theta)$ in $H^{1}_{per}(\square;\mathbb{C}^{2})$, with coercivity constant $\nu(\theta)>0$. In fact, we can say much more; the form $a_{\theta}$ is uniformly coercive on $(\mathcal{V}(\theta))^{\perp}$, i.e. $\nu_{*}:=\inf_{\theta}\nu(\theta)>0$. Indeed, it was shown in [5, Theorem 3.4] that the uniform coercivity of $a_{\theta}$ on the orthogonal complement to its kernel is equivalent to the kernel being continuous with respect to $\theta$. We demonstrated in Theorem 4.1 that $V(\theta)$ was Lipschitz continuous in $\theta$, thus clearly $\mathcal{V}(\theta)$ is Lipschitz in $\theta$ and consequently Theorem 3.2 of [5] holds, which, in this setting, implies:
@@
$\left|b_{\varepsilon}(\phi,\psi)-b(\phi,\psi)\right|\leq\varepsilon L_{0}\|\nabla\phi\|_{L^{2}(\square;\mathbb{C}^{2})}\|\nabla\psi\|_{L^{2}(\square;\mathbb{C}^{2})}\quad\forall\phi,\psi\in H^{1}(\square;\mathbb{C}^{2}),$ (6.11)
@@
$|\sqrt{1-\varepsilon}-1|\leq\varepsilon,\quad\text{and}\quad|\sqrt{1-\varepsilon}-(1-\tfrac{\varepsilon}{2})|\leq\frac{\varepsilon^{2}}{2},\quad\varepsilon\in(0,1),$
@@
$\left|b_{\varepsilon}(\phi,\psi)-b(\phi,\psi)\right|\leq\frac{\varepsilon}{2}\int_{\square\backslash R}|\nabla\phi_{1}^{\perp}\cdot\overline{\nabla\psi_{2}}+\nabla\phi_{2}\cdot\overline{\nabla^{\perp}\psi_{1}}|+$
@@
$-\Delta u_{i}=\chi_{R_{\delta}}f_{i}\quad\text{in $\square$}.$
@@
$-\Delta v_{i}=\chi_{R_{\delta}}f_{i}\quad\text{in $\mathbb{R}^{2}$ with $v_{i}(x)=-\frac{d_{i}}{2\pi}\ln{|x|}+o(1)$ as $|x|\rightarrow+\infty$.}$
@@
${b}^{(\delta)}(f,\tilde{f}):=\gamma^{-1}\int_{R_{\delta}}f\cdot\overline{\tilde{f}}-\int_{R_{\delta}}\left[\Big{(}(A_{\delta}f_{1})_{,22}+(A_{\delta}f_{2})_{,12}\Big{)}\overline{\tilde{f}_{1}}+\Big{(}(A_{\delta}f_{1})_{,21}+(A_{\delta}f_{2})_{,11}\Big{)}\overline{\tilde{f}_{2}}\right],$ (6.12)
@@
$\left|\frac{b_{\theta}[f]}{c_{\theta}[f]}-\frac{b^{(\delta)}[f]}{c_{\delta,\theta}[f]}\right|\leq C_{1}\delta\left(\frac{b_{\theta}[f]}{c_{\theta}[f]}+\frac{b^{(\delta)}[f]}{c_{\delta,\theta}[f]}\right),\qquad\forall f\in L^{2}(R_{\delta};\mathbb{C}^{2}),$
@@
[…]
@@
$\langle\partial_{r}v_{i},h\rangle_{S_{\beta}}$ $=-\langle v_{i},\partial_{r}h\rangle_{S_{\beta}}-d_{i}\overline{h(0)};\quad\text{for $\delta\leq\beta\leq\alpha$;}$ (6.16)
@@
$v_{i}(r,\theta)=-\frac{d_{i}}{2\pi}\ln{r}+\sum_{n=1}^{\infty}c_{n}r^{-n}e^{\mathrm{i}n\theta},\qquad\text{in $\mathbb{R}^{2}\backslash B_{\delta}$},$ (6.19)
@@
$h(r,\theta)=h(0)+\sum_{n=1}^{\infty}h_{n}r^{n}e^{\mathrm{i}n\theta},\qquad\text{in $B_{\alpha}$},$ (6.20)
@@
$\langle\nabla h,\nabla h\rangle_{B_{\beta}}=\langle\partial_{r}h,h\rangle_{S_{\beta}}=2\pi\sum_{n=1}^{\infty}n\beta^{2n}|h_{n}|^{2}\qquad\text{for $\delta\leq\beta\leq\alpha$.}$
@@
00 0 $=\langle G_{\theta},-\Delta G_{\theta}\rangle_{\square\backslash B_{\alpha}}=\|\nabla G_{\theta}\|_{\square\backslash B_{\alpha}}^{2}+\langle G_{\theta},\partial_{r}G_{\theta}\rangle_{S_{\alpha}},$
@@
$\langle g_{\theta},\partial_{r}2\pi^{-1}\ln{r}\rangle_{S_{\alpha}}=g_{\theta}(0)\quad\text{and}\quad\langle 2\pi^{-1}\ln{r},\partial_{r}g_{\theta}\rangle_{S_{\alpha}}=0.$
@@
=(6.16)${{=}}\langle\nabla v_{i},\nabla w_{j}\rangle_{B_{\delta}}+\langle v_{i},\partial_{r}w_{j}\rangle_{S_{\delta}}+d_{i}\overline{w_{j}(0)}$
@@
=(6.16)${{=}}\langle\nabla v_{j},\nabla g_{\theta}\rangle_{B_{\delta}}+\langle v_{j},\partial_{r}g_{\theta}\rangle_{S_{\delta}}+d_{j}\overline{g_{\theta}(0)}$
@@
(6.15)${{\leq}}2\frac{\delta}{\alpha}\Big{(}\|\nabla v_{i}\|_{B_{\alpha}}\|\nabla w_{j}\|_{B_{\alpha}}+|d_{i}|\|\nabla g_{\theta}\|_{B_{\alpha}}\|\nabla v_{j}\|_{B_{\alpha}}\Big{)}$ )
@@
00 0 $=\langle-\Delta u_{j},u_{j}\rangle_{\square\backslash B_{\alpha}}=\|\nabla u_{j}\|^{2}_{\square\backslash B_{\alpha}}+\langle\partial_{r}u_{j},u_{j}\rangle_{S_{\alpha}}$
@@
=(6.21)${{=}}\langle f_{j},v_{j}\rangle+|d_{j}|^{2}g_{\theta}(0)+\mathrm{Re}(2d_{j}\langle\nabla g_{\theta},\nabla v_{j}\rangle_{B_{\delta}})$ )
@@
(6.15)${{\leq}}\frac{\delta}{\alpha}\big{(}\|v_{i,j}\|_{B_{\alpha}}\|w_{k,l}\|_{B_{\alpha}}+\|w_{i,j}\|_{B_{\alpha}}\|u_{k,l}\|_{B_{\alpha}}\big{)}$ )
@@
$\langle g_{,i},(A_{\delta}g)_{,i}\rangle\leq\langle g,g\rangle,\quad g\in C^{\infty}_{0}(R_{\delta}),\ i=1,2.$

### v2

@@ 6 Appendix
$(u,\tilde{u})_{H^{1}_{\theta}(\square;\mathbb{C}^{2})}=\left\{\begin{array}[]{lcc}\int_{\square}\nabla u:\overline{\nabla\tilde{u}}&&\theta\neq 0,\\[1.7pt] \int_{\square}\nabla u:\overline{\nabla\tilde{u}}+\left|\int_{\square}u\right|\left|\int_{\square}\tilde{u}\right|&&\theta=0,\end{array}\right$. (6.1)
@@
$\text{$\mathrm{div}\,{w}=0\qquad$ and $\qquad\mathrm{curl}\,{w}=0\qquad$ in $C$},$
@@
[…]
@@
$\qquad\forall\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}),$
@@
$\int_{\square\backslash R}c_{1}\overline{\mathrm{div}\,{\phi}}+c_{2}\overline{\mathrm{curl}\,{\phi}}+\int_{\square\backslash R}v_{2,1}\overline{\left(\frac{\phi_{2,1}+\phi_{1,2}}{2}\right)}+v_{2,2}\overline{\left(\frac{\phi_{2,2}-\phi_{1,1}}{2}\right)}$
@@
[…]
$\int_{\square\backslash R}c_{1}$ $\overline{\mathrm{div}\,{\tilde{w}}}+c_{2}\overline{\mathrm{curl}\,{\tilde{w}}}=\langle F,\tilde{w}\rangle,\qquad\forall\tilde{w}\in W(\theta),$ (6.4)
@@
[…]
$-\int_{\square\backslash R}v_{2,1}\overline{\left(\frac{\tilde{w}_{2,1}+\tilde{w}_{1,2}}{2}\right)}+v_{2,2}\overline{\left(\frac{\tilde{w}_{2,2}-\tilde{w}_{1,1}}{2}\right)}-\int_{R}\nabla v_{2}\cdot\overline{\nabla\tilde{w}_{2}}.$
@@
$\int_{\square\backslash R}\mathrm{div}\,{w}\,\overline{\mathrm{div}\,\tilde{w}}+\mathrm{curl}\,{w}\,\overline{\mathrm{curl}\,{w}}=\langle F,\tilde{w}\rangle\qquad\forall\tilde{w}\in W(\theta).$ (6.5)
@@
Here, we shall prove Lemma 3.1. For this, we shall rely on the techniques introduced in [8, Section 3].
For fixed $0<\varepsilon<1$, $\theta\in\square^{*}$ and $F\in H^{-1}(\square;\mathbb{C}^{2})$, recall $U_{\varepsilon,\theta}$ is the solution to
$\varepsilon^{-1}a(U_{\varepsilon,\theta},\phi)+b_{\varepsilon}(U_{\varepsilon,\theta},\phi)+c(U_{\varepsilon,\theta},\phi)=\langle F,\phi\rangle,\qquad\forall\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}),$ (6.6)
@@
$b(V_{\theta},\tilde{V})=\langle F,\tilde{V}\rangle,\qquad\forall\,\tilde{V}\in V(\theta).$
@@
$\varepsilon^{-1}a(W_{\varepsilon,\theta},\phi)+b(W_{\varepsilon,\theta},\phi)+c(W_{\varepsilon,\theta},\phi)=\langle F,\phi\rangle,\qquad\forall\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}),$ (6.7)
@@
$\nu_{1}\|\nabla\phi\|^{2}_{L^{2}(\square;\mathbb{C}^{2})}\leq a[\phi]+b[\phi]\leq\nu_{2}\|\nabla\phi\|^{2}_{L^{2}(\square;\mathbb{C}^{2})},\qquad\forall\phi\in H^{1}(\square;\mathbb{C}^{2}),$ (6.8)
@@
$\left|b_{\varepsilon}(\phi,\psi)-b(\phi,\psi)\right|\leq\varepsilon L_{b}\sqrt{\varepsilon^{-1}a[\phi]+b_{\varepsilon}[\phi]}\sqrt{a[\psi]+b[\psi]}\qquad\forall\phi,\psi\in H^{1}(\square;\mathbb{C}^{2}),$ (6.9)
and for all $0<\varepsilon<1$.
@@
for some $\nu_{*}$ independent of $\varepsilon,\theta$ and $F$. To do this, we shall utilise [8, Theorem 3.1]; setting,
$a_{\theta}(\phi,\psi):=a(e^{\mathrm{i}\theta\cdot y}\phi,e^{\mathrm{i}\theta\cdot y}\psi),\hskip 17.00024pt\text{and}\qquad\ b_{\theta}(\phi,\psi):=b(e^{\mathrm{i}\theta\cdot y}\phi,e^{\mathrm{i}\theta\cdot y}\psi)+c(\phi,\psi),$
@@
$\varepsilon^{-1}a_{\theta}(u_{\varepsilon,\theta},\phi)+b_{\theta}(u_{\varepsilon,\theta},\phi)=\langle f,\phi\rangle,\qquad\forall\phi\in H^{1}_{per}(\square;\mathbb{C}^{2});$
@@
$b_{\theta}(v_{\theta},\tilde{v})=\langle f,\tilde{v}\rangle,\qquad\forall\tilde{v}\in\mathcal{V}(\theta).$
@@
$\mathcal{V}(\theta):=\{v\in H^{1}_{per}(\square;\mathbb{C}^{2})\,|\,\mathrm{div}\,{v}+\mathrm{i}\theta\cdot v=0,\text{ and }\mathrm{curl}\,{v}+\mathrm{i}\theta^{\perp}\cdot v=0\text{ in $\square\backslash R$}\}.$
We now argue as in [8, Section 3]. First note that the closed subspace $\mathcal{V}(\theta)$ is the kernel of the form $a_{\theta}$, precisely
@@
Furthermore, it can be seen that $a_{\theta}$ is coercive on $(\mathcal{V}(\theta))^{\perp}$, the orthogonal complement of $\mathcal{V}(\theta)$ in $H^{1}_{per}(\square;\mathbb{C}^{2})$, with coercivity constant $\nu(\theta)>0$. In fact, we can say much more; the form $a_{\theta}$ is uniformly coercive on $(\mathcal{V}(\theta))^{\perp}$, i.e. $\nu_{*}:=\inf_{\theta}\nu(\theta)>0$. Indeed, it was shown in [8, Theorem 3.4] that the uniform coercivity of $a_{\theta}$ on the orthogonal complement to its kernel is equivalent to the kernel being continuous with respect to $\theta$. We demonstrated in Theorem 4.1 that $V(\theta)$ was Lipschitz continuous in $\theta$, thus clearly $\mathcal{V}(\theta)$ is Lipschitz in $\theta$ and consequently Theorem 3.2 of [8] holds, which, in this setting, implies:
@@
$\left|b_{\varepsilon}(\phi,\psi)-b(\phi,\psi)\right|\leq\varepsilon L_{0}\|\nabla\phi\|_{L^{2}(\square;\mathbb{C}^{2})}\|\nabla\psi\|_{L^{2}(\square;\mathbb{C}^{2})}\qquad\forall\phi,\psi\in H^{1}(\square;\mathbb{C}^{2}),$ (6.11)
@@
$|\sqrt{1-\varepsilon}-1|\leq\varepsilon,\qquad\text{and}\qquad|\sqrt{1-\varepsilon}-(1-\tfrac{\varepsilon}{2})|\leq\frac{\varepsilon^{2}}{2},\qquad\varepsilon\in(0,1),$
@@
[…]
@@
$\Delta u_{i}=\chi_{R_{\delta}}f_{i}\qquad\text{in $\square$}.$
@@
$-\Delta v_{i}=\chi_{R_{\delta}}f_{i}\qquad\text{in $\mathbb{R}^{2}$ with $v_{i}(x)=-\frac{d_{i}}{2\pi}\ln{|x|}+o(1)$ as $|x|\rightarrow+\infty$.}$
@@
${b}^{(\delta)}(f,\tilde{f})=\gamma^{-1}\int_{R_{\delta}}f\cdot\overline{\tilde{f}}-\int_{R_{\delta}}\left[\Big((A_{\delta}f_{1})_{,22}+(A_{\delta}f_{2})_{,12}\Big)\overline{\tilde{f}_{1}}+\Big((A_{\delta}f_{1})_{,21}+(A_{\delta}f_{2})_{,11}\Big)\overline{\tilde{f}_{2}}\right],$ (6.12)
@@
$\left|\frac{b_{\theta}[f]}{c_{\theta}[f]}-\frac{b^{(\delta)}[f]}{c_{\delta,\theta}[f]}\right|\leq C_{1}\delta\left(\frac{b_{\theta}[f]}{c_{\theta}[f]}+\frac{b^{(\delta)}[f]}{c_{\delta,\theta}[f]}\right),\hskip 17.00024pt\forall f\in L^{2}(R_{\delta};\mathbb{C}^{2}),$
@@
For ease of exposition we shall use, henceforth, $\langle\cdot,\cdot\rangle$ to denote $\langle\cdot,\cdot\rangle_{R_{\delta}}$. Let us prove Lemma 6.2. The following, key estimate, is instrumental to the proof:
@@
$\langle\partial_{r}v_{i},h\rangle_{S_{\beta}}$ $=-\langle v_{i},\partial_{r}h\rangle_{S_{\beta}}-d_{i}\overline{h(0)};\qquad\text{for $\delta\leq\beta\leq\alpha$;}$ (6.16)
@@
[…]
@@
$h(r,\theta)=h(0)+\sum_{n=1}^{\infty}h_{n}r^{n}e^{\mathrm{i}n\theta},\hskip 17.00024pt\text{in $B_{\alpha}$},$ (6.20)
@@
$\langle\nabla h,\nabla h\rangle_{B_{\beta}}=\langle\partial_{r}h,h\rangle_{S_{\beta}}=2\pi\sum_{n=1}^{\infty}n\beta^{2n}|h_{n}|^{2}\hskip 17.00024pt\text{for $\delta\leq\beta\leq\alpha$.}$
@@
$0$ $=\langle G_{\theta},-\Delta G_{\theta}\rangle_{\square\backslash B_{\alpha}}=\|\nabla G_{\theta}\|_{\square\backslash B_{\alpha}}^{2}+\langle G_{\theta},\partial_{r}G_{\theta}\rangle_{S_{\alpha}},$
@@
$\langle g_{\theta},\partial_{r}2\pi^{-1}\ln{r}\rangle_{S_{\alpha}}=g_{\theta}(0)\qquad\text{and}\qquad\langle 2\pi^{-1}\ln{r},\partial_{r}g_{\theta}\rangle_{S_{\alpha}}=0.$
@@
[…]
@@
[…]
@@
[…]
@@
$0$ $=\langle-\Delta u_{j},u_{j}\rangle_{\square\backslash B_{\alpha}}=\|\nabla u_{j}\|^{2}_{\square\backslash B_{\alpha}}+\langle\partial_{r}u_{j},u_{j}\rangle_{S_{\alpha}}$
@@
[…]
@@
[…]
@@
$\langle g_{,i},(A_{\delta}g)_{,i}\rangle\leq\langle g,g\rangle,\qquad g\in C^{\infty}_{0}(R_{\delta}),\ i=1,2.$